# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

### v2

@@ 5 Retrieval analysis
After detecting the species shown in Sect. 3.1 thanks to cross-correlation analysis, the next step is to explore the abundances of these species in comparison to solar values (This comparison is possible due to the solar metallicity of the host star). One approach developed in Brogi & Line 2019 is to use a Bayesian atmospheric retrieval framework with high-resolution cross-correlation spectroscopy (HRCCS) that relies on the cross-correlation between data and models for extracting the planetary spectral signal. This approach permits the characterization of many atmospheres of UHJs and puts constraints on abundances (Line et al., 2021; Kasper et al., 2021, 2023; Brogi et al., 2023).
5.1 CHIMERA codes
Following the above method, we applied the Brogi & Line 2019 cross-correlation-to-log-likelihood retrieval framework to derive the molecular volume mixing ratios and the temperature layer we are probing. For the retrieval process, we used the basic CHIMERA “free-retrieval” (Line et al., 2013; Kreidberg et al., 2015) paradigm, which assumes constant-with-altitude gas mixing ratios and uses a simple isothermal T-P profile. Pelletier et al. 2023 shows that a simple isothermal profile would give a similar estimation of chemical abundance ratio to a more complex T-P profile. The retrieval parameters specific to our analysis and their prior ranges are provided in Table 5. A more detailed description of the high-resolution GPU-based radiative transfer method and loglikelihood implementation within pymultinest (Feroz et al., 2009; Buchner et al., 2014) is given in Line et al. 2021. Retrieval parameters and their priors are reported in Table 5.
@@ 5.1 CHIMERA codes
| Log(a) | Model scaling factor | -2 - 2 |
| Log(H-) | Continuum | -12 - 0 |
| Log(Fe) | log gas volume mixing ratio | -12 - 0 |
| Log(/Fe) | Relative abundance | -8 - 1 |
| Log() | Ionised fraction (for Fe,Ti,Ca) | -2 - 0 |
@@ 5.2 Results of the retrieval analysis
We combined the blue and red arm MAROON-X data for our retrieval for both transits. We included most of the species detected in Sect. 3.1 TiO and abundance proxies for the H bound-free and free-free continua. For detected species in both ionized and neutral states, we compute an ionized fraction of them. (Example with Fe in equation 2):
Because we only have a 3-sigma upper limit of 2.3 M for the planetary mass (Cabot et al., 2021), we include the mass as a free parameter that spanned a range between maximum Mp = 2.3 M and Mp = 0.5 M. For this analysis, we consider the two detectors independently as for the CCF, and we add the likelihood from the red and the blue detector analysis and then from both transit. For the retrieval analysis, we used three number of principal components as for the CCF analysis. Because we had to put the same number of principal component for each species, the Ca+ lines may be affected a bit by the PCA. The results are summarized in Fig. 14 and Fig. 15 and the whole corner plot is given in the Appendix (Fig. 24).
@@ 5.3 Discussion of the retrieval analysis
The and Vres distribution of the retrieval analysis (Fig. 14) are consistent with the iron signal detected in section 3.1. The measured Vres () and () also agree with the detection of Cabot et al. (2021) ( km/s and = km/s). We observed that the velocity parameters obtained are influenced predominantly by the strongest absorber, particularly Fe. Other species may have minor differences in and , as discussed in section 3.1. However, we have chosen not to individually adjust and for each species to derive meaningful abundance ratios. Some species detected in the Cross-Correlation analysis were not included in the retrieval framework due to a lack of opacities files like Ba+, Si, Mn, and the HyVO line list for VO (for which we used the VOmyt line list instead).
Our retrievals converge towards values of Mg, Ca, and Fe abundances on TOI-1518 b that are consistent with proto-solar abundances. Cr, Ti, and V, however, differ at the 1 level. The lower abundances of Cr, V and Ti can plausibly be explained by their condensation on the planet nightside or by the fact that they are bound to other molecules (e.g. TiO, VO), or by the fact that they are partially ionized ( e.g. Na+, Cr+). We already showed that the line lists are essential for the cross-correlation (Fig. 8). We could not use the HyVO line list in the retrieval analysis. The VOmyt line list leads to a nondetection of VO in the retrieval analysis, as in the CCF analysis. The missing vanadium is then probably present in VO or is ionized. For titanium, the subsolar value could also be explained by the non-detection of TiO. Further observations with bluer or redder instruments to detect molecules or ionized species are needed to determine why such elements are subsolar.
Even if the isothermal TP profile could bias our results, Pelletier et al. 2023 showed that different temperature-pressure profiles inferred similar elemental abundance ratios.
@@ 6 Conclusions
This study presents an in-depth analysis of TOI-1518 b, an ultra-hot Jupiter. It uses transit observations, from MAROON-X on Gemini-N, to explore its atmospheric dynamics and chemical composition. Our findings offer new insights into the unique characteristics of this extreme exoplanet.
Our cross-correlation analysis focused on detecting atomic and molecular species within the atmosphere. We report the detection of 14 different species. High-resolution spectroscopy allowed us to identify ionized and neutral species with ionized metals such as Fe+, Ca+, and Ti+. The detection of these ionized species, as opposed to their neutral counterparts, underscores the extreme temperatures of TOI-1518 b’s atmosphere, where thermal ionization is significant. Additionally, we reported the detection of vanadium oxide (VO), a critical absorber that has implications for thermal inversions in ultra-hot Jupiters. The presence of VO adds a crucial piece to the puzzle of understanding the thermal structure and chemical processes in such extreme environments. This was possible thanks to the newer HyVO line list that sould be used in other studies were detection of VO failed with previous line lists
We investigated the atmospheric wind dynamics by analyzing the blueshift of multiple species in the observed spectra as a function of the planet’s orbital phase. The blueshift of iron is consistent with previous observations of other ultra-hot Jupiters. By combining the signal from different species, particularly Fe+ and Fe, and comparing them with Global Circulation Models, we conclude that a strong drag is needed (. Furthermore, ionized species need more substantial drags compared to neutral species, probably because of the amplified magnetic effects present in the upper atmosphere.
The retrieval analysis provided constraints on the abundance of various chemical species in the atmosphere. Our results suggest non-solar elemental abundance ratios for Vanadium, Magnesium, Chromium, and Titanium relative to iron, highlighting the complex chemistry of TOI-1518 b.
Overall, this study demonstrates the power of combining high-resolution spectroscopy with advanced modeling techniques to probe the atmospheres of ultra-hot Jupiters. They shed light on the specific properties of TOI-1518 b and contribute to the broader understanding of atmospheric dynamics and chemistry in these extreme exoplanets. Future studies should continue to refine these models and expand observational efforts to explore the diversity and complexity of ultra-hot Jupiter atmospheres.
@@ Acknowledgements.
This project has received funding from Agence Nationale de la Recherche (ANR) for the grant ANR EXOWINDS.This work was supported by the French government through the France 2030 investment plan managed by the National Research Agency (ANR), as part of the Initiative of Excellence Université Côte d’Azur under reference number ANR-15-IDEX-01. The authors are grateful to the Université Côte d’Azur’s Center for High-Performance Computing (OPAL infrastructure) for providing resources and support.
@@
Appendix A Cross-Correlation-Function
Appendix B Global Circulation Models
Appendix C Retrieval analysis

### v3

@@ 5 Retrieval analysis
After detecting the species shown in Sect. 3.2 thanks to cross-correlation analysis, the next step is to explore the abundances of these species in comparison to solar values (This comparison is possible due to the solar metallicity of the host star). One approach developed in Brogi & Line 2019 is to use a Bayesian atmospheric retrieval framework with high-resolution cross-correlation spectroscopy (HRCCS) that relies on the cross-correlation between data and models for extracting the planetary spectral signal. This approach allows the characterization of many atmospheres of UHJs and puts constraints on abundances (Line et al., 2021; Kasper et al., 2021, 2023; Brogi et al., 2023).
5.1 CHIMERA code
Following the above method, we applied the Brogi & Line 2019 cross-correlation-to-log-likelihood retrieval framework to derive the molecular volume mixing ratios and the temperature layer we are probing. For the retrieval process, we used the CHIMERA “free-retrieval” (Line et al., 2013; Kreidberg et al., 2015) paradigm, which assumes constant-with-altitude gas mixing ratios and uses a simple isothermal T-P profile, which is a good enough approximation to retrieve abundances with MAROON-X as shown by Pelletier et al. 2023. We use a similar setup as Line et al. 2021 but with a wavelength range and a choice of chemical species adapted to the MAROON-X bandpass. Additionally, we use a native resolution for the radiative transfer of 500.000, that is downgraded to the MAROON-X 85.000 resolution after instrumental and rotational broadening and orbital Doppler shift. The retrieval parameters specific to our analysis and their prior ranges are provided in Table 5. A more detailed description of the high-resolution GPU-based radiative transfer method and log-likelihood implementation within pymultinest (Feroz et al., 2009; Buchner et al., 2014) is given in Line et al. 2021.
@@ 5.1 CHIMERA code
| a) | Model scaling factor | -2 - 2 |
| H-) | Continuum | -12 - 0 |
| ) | log gas volume mixing ratio | -12 - -1.5 |
@@ 5.2 Results of the retrieval analysis
We combined the blue and red arm MAROON-X data for our retrieval for both transits. We included most of the species detected in Sect. 3.2 TiO and abundance proxies for the H bound-free and free-free continua. However, some species detected in the Cross-Correlation analysis were not included in the retrieval framework due to a lack of opacities files in the right format, like Ba+, Si and Mn.
The mass prior is set using the SOPHIE observation results (see Table 7) as a free parameter that follows a Gaussian distribution with Mp = 1.83 M. For this analysis, we consider the two detectors independently as for the CCF, and we add the likelihood from the red and the blue detector analysis and then from both transit. For the retrieval analysis, we used three number of principal components as for the CCF analysis. As shown before, the PCA does not strongly affect the planetary lines, apart from the ones of the . However, because the retrieval considers all species at the same time, we could not have different PCA numbers per species. The results of the measured abundances are summarized in Fig. 14 and the whole corner plot is given in the Appendix.(Fig. 26).
@@ 5.3 Discussion of the retrieval analysis
The Kp and Vres distribution of the retrieval analysis(Fig. 26) are consistent with the iron signal detected in section 3.2. The measured Vres () and Kp () also agree with the detection of Cabot et al. (2021) ( km/s and Kp= km/s). We observed that the velocity parameters obtained are influenced predominantly by the strongest absorber, particularly Fe.
Our model retrieves an iron abundance of . This corresponds to 0.07 to 1.62 solar enrichment in iron. This remains consistent with a solar (and then stellar) metallicity. We have checked that the iron abundance was robust to different assumptions in the retrieval, and that it was not changing when different other species were added.
The retrieval results for the other chemical species are more surprising and highlight the difficulty of atmospheric retrievals for transit spectroscopy at high-spectral resolution. First, the ionized species, such as and have retrieved abundances that are much higher than solar. These high abundances, up to are unlikely in a gas giant atmosphere. Instead, we believe that these could stem from a lack of flexibility in the forward model. Indeed, ionized species are likely to probe different parts of the atmosphere (both vertically and longitudinally) than neutral species, because the spatial repartition of neutral and ionized forms of the same species is anti-correlated. This can lead to signals at different temperatures and Kp and Vres (e.g. Table 3). However, given that the temperature, Kp and Vres for all species is determined by the neutral iron lines, which represent the strongest signal, the abundances of the other species are likely biased to compensate for the Kp/Vres offset.
For lines with a strong Kp/Vsys shift and small opacities, such as , , or , the retrieval leads only to upper limits, which are all consistent with a solar enrichment.
For the neutral calcium and magnesium lines, the retrieval provides a detection with an abundance of and , corresponding to 0.0002 to 0.034 times the solar abundance for and 0.0004 to 0.339 times solar for . The retrieved value for neutral calcium is significantly lower than that of iron, and it is, for now, unclear whether this is due to a physical effect (e.g. most being into ) or a bias in the retrieval. Finally, our results for deserve a special mention.
Despite using the latest line list, the multi-species retrieval did not successfully detect . To explore this issue further, we conducted a single-species retrieval focused solely on (see Fig. 27). This approach resulted in a detection with a temperature around 2700 K, compared to 3500 K in the multi-species retrieval, along with a different distribution (Fig. 15). The measured value of in such a framework is , corresponding to an enrichment of 1.10 to 45708.81 times solar. The discrepancy in the distribution resembles the patterns observed in the map (see Table 4). This difference could be attributed to either dynamical factors—if probes specific regions of the atmosphere—or inaccuracies in the positioning of the lines within the line list. This situation underscores the limitations of the current framework in accurately deriving abundance for species that are settled in different distribution.
Further observations with bluer or redder instruments to detect molecules or ionized species are needed to determine why such elements are subsolar. Moreover, using a non-isothermal profile instead would have a dual effect on the spectrum. First, it would produce a similar effect to an abundance gradient by stretching the spectral lines. Second, it would enhance the abundance gradient in the case of chemical equilibrium, further impacting the observed spectral features. Incorporating more realistic temperature and abundance gradients in future retrieval analyses will be essential to better constrain the abundances of ionized species.
@@ 6 Conclusions
This study presents an in-depth analysis of TOI-1518 b, an ultra-hot Jupiter. It uses transit observations, from MAROON-X on Gemini-N, to explore its atmospheric dynamics and chemical composition. Our findings offer new insights into the unique characteristics of this extreme exoplanet. Using the SOPHIE spectrograph, we also report the first significant detection of TOI-1518 b in radial velocity. This allows us to better characterize the system, and in particular to measure the planetary mass MJup.
Our cross-correlation analysis focused on detecting atomic and molecular species within the atmosphere. We report the detection of 14 different species. High-resolution spectroscopy allowed us to identify ionized and neutral species with ionized metals such as , , and . The detection of these ionized species, as opposed to their neutral counterparts, underscores the extreme temperatures of TOI-1518 b’s atmosphere, where thermal ionization is significant. Additionally, we reported the detection of vanadium oxide (), a critical absorber that has implications for thermal inversions in ultra-hot Jupiters. The presence of VO adds a crucial piece to the puzzle of understanding the thermal structure and chemical processes in such extreme environments. This was possible thanks to the newer HyVO line list that should be used in other studies were detection of VO failed with previous line lists.
We investigated the atmospheric wind dynamics by analyzing the blueshift of multiple species in the observed spectra as a function of the planet’s orbital phase. The blueshift of iron is consistent with previous observations of other ultra-hot Jupiters. By combining the signal from different species, particularly and Fe, and comparing them with global circulation models , we conclude that a strong drag is needed ( to explain the data. seem to need more substantial drag compared to Fe, which can be due to the increased Ohmic drag in the planetary dayside. Finally, we find that the trail of is very different from the other species, which is likely due to probing the escaping atmosphere above the Roche lobe while all other species probe deeper atmospheric pressures.
The retrieval analysis provided constraints on the abundance of various chemical species in the atmosphere. We report an abundance of iron of , which is consistent with a solar enrichment. The retrieval being driven mainly by iron, we believe that the iron abundances are robust. However, most other species are detected at a different /, which our retrieval framework does not consider. As a consequence, we believe that the retrieved abundance for the other species are strongly biased. We study in more detail this bias for the specific case of , for which our main retrieval does not find a signal, but where a retrieval without iron lead to a detection at a different temperature, , and an abundance of 1.10 to 45708.81 times solar.
Overall, our study demonstrates the power of combining high-resolution spectroscopy with advanced modeling techniques to probe the atmospheres of ultra-hot Jupiters. They shed light on the specific properties of TOI-1518 b and contribute to the broader understanding of atmospheric dynamics and chemistry in these extreme exoplanets. Future studies should continue to refine these models and expand observational efforts to explore the diversity and complexity of ultra-hot Jupiter atmospheres.
@@ Acknowledgements.
This work was partially funded by the French National Research Agency (ANR) project EXOWINDS (ANR-23-CE31-0001-01).This work was supported by the French government through the France 2030 investment plan managed by the National Research Agency (ANR), as part of the Initiative of Excellence Université Côte d’Azur under reference number ANR-15-IDEX-01. The authors are grateful to the Université Côte d’Azur’s Center for High-Performance Computing (OPAL infrastructure) for providing resources and support.
@@
This work uses observations secured with the SOPHIE spectrograph at the 1.93-m telescope of Observatoire Haute-Provence, France, with the support of its staff.
This work was supported by the “Programme National de Planétologie” (PNP) of CNRS/INSU, and CNES.
Appendix A Measurement of the planetary mass
Cabot et al. 2021 obtained a hint (less than 2 significance) for a radial-velocity (RV) detection of TO-1518b using the FIES spectrograph. They reported an RV semi-amplitude of m/s, which they converted to a 2- upper limit of 281 m/s, corresponding to a planetary upper mass limit of 2.3 MJup.
With the goal to improve the constraint on the planetary mass, we observed TOI-1518 with the SOPHIE spectrograph at the 1.93-m telescope of the Observatoire de Haute-Provence, France. SOPHIE is a stabilized échelle spectrograph dedicated to high-precision RV measurements (Perruchot et al. 2008; Bouchy et al. 2009, 2013). We used its high-resolution mode (resolving power ) and fast readout mode. We obtained exposures at 14 different epochs between December 2023 and January 2024. Exposure times ranged between 3.9 and 16.8 minutes depending on weather conditions, allowing signal-to-noise from 35.6 to 47.3 to be reached per pixel at 550 nm. The log of the observations is reported in Table 6.
The RVs were extracted with the SOPHIE pipeline, as presented by Bouchy et al. (2009) and refined by Heidari et al. (2024). It derives cross correlation functions (CCF) from standard numerical masks corresponding to different spectral types. As expected from the known rapid rotation speed of the star, the derived CCFs were broad, with typical FWHM around 100 km/s. So to measure the RV, instead of fitting a Gaussian profile as it is standardly done for slow rotators, here we fitted the CCF with a profile constructed from the convolution of a Gaussian and a rotational profile following Gray (2022). This provided good fits of the CCFs. The Moonlight pollution, estimated using the second SOPHIE fiber aperture that is targeted on the sky, 2’ away from the first one pointing toward the star, was negligible.
The FWHM of the Gaussian is unknown, but is expected to be small by comparison to the FWHM of the rotational profile. We attempted free and fixed values for that Gaussian FWHM, which did not significantly modify our results. We finally fixed the FWHM of the Gaussian profile to 7 km/s on each exposure, which is a common value for such kind of stars. The FWHM of the rotational profil was free to vary among exposures, and we obtained values corresponding to km/s, with a dispersion of km/s. This is in good agreement with the values km/s reported by Cabot et al. 2021. The measured RVs (see Table 6) are derived from the fitted center of the Gaussian and rotational profiles, constrained to be identical for a given epoch.
We phase-folded those SOPHIE’s RVs using the 1.9-d period derived from the transits observed by TESS. This provided a significant RV semi-amplitude m/s, in phase with the transit epoch measured with TESS. This agrees with the upper limit reported by Cabot et al. (2021). So we concluded our SOPHIE data allow the planet TOI-1518 b to be significantly detected, for the first time with RVs.
To refine the system parameters, we made a joined fit of the TESS transit light curves and the available RVs. Following Heidari et al. (2025), we fitted that dataset using the EXOFASTv2 package (Eastman et al. 2013; Eastman 2017; Eastman et al. 2019). Cabot et al. (2021) used TESS sectors 17 and 18 covering TOI-1518, observed in FFI mode so with a long-cadence 30-min sampling, in addition to five ground-based photometric follow up. Here, we used both those TESS sectors, as well as the three additional TESS sectors covering TOI-1518 now available, which have a short-cadence 2-min sampling (sectors 57, 58, and 77). So by comparison to Cabot et al. (2021), the number of TOI-1518 b individual transits included in our fit increases from 24 to 56. And in addition to our new SOPHIE RVs, we also used the FIES RVs published by Cabot et al. (2021).
Cabot et al. (2021) reported a negligible eccentricity (). We tested eccentric and circular orbits without detecting significant differences, and finally adopted here a circular orbit. According to Watanabe et al. (2024), the system shows a change in the impact parameter of the planet as a function of time. Over the 4.5-yr span of the TESS observations we use, the expected effect is of the order of 0.05 on . This is not taken into account here, and we report averaged values. This has no significant impact on the reported planetary mass.
Our fitted RVs and transit light curve are shown in Fig. 16 and 17, respectively. We finally measured an RV semi-amplitude m/s, corresponding to a planetary mass MJup. The full results of our fit are reported in Table 7. Our derived parameters agree with those presented by Cabot et al. 2021, with improved uncertainties for most of them.
| BJDUTC | RV | exp. | SNR per pix | |
| -2 460 000 | (km/s) | (km/s) | (sec) | at 550 nm |
| 287.28150 | -11.64 | 0.11 | 1010 | 46.3 |
| 294.30716 | -11.89 | 0.14 | 543 | 46.7 |
| 295.25968 | -11.30 | 0.13 | 284 | 46.6 |
| 296.31131 | -11.68 | 0.12 | 268 | 46.9 |
| 297.29585 | -11.46 | 0.12 | 271 | 46.6 |
| 298.23770 | -11.60 | 0.12 | 233 | 46.5 |
| 300.28312 | -12.12 | 0.13 | 803 | 41.6 |
| 302.22014 | -12.21 | 0.13 | 662 | 46.4 |
| 323.26383 | -11.88 | 0.13 | 452 | 47.1 |
| 334.29677 | -11.92 | 0.16 | 904 | 35.6 |
| 335.24376 | -11.60 | 0.12 | 593 | 46.6 |
| 336.24214 | -12.10 | 0.12 | 530 | 46.9 |
| 337.23923 | -11.52 | 0.12 | 237 | 46.7 |
| 339.25846 | -11.46 | 0.11 | 407 | 47.3 |
| Parameter | Units | Values | |
| Stellar Parameters: | TOI-1518 | ||
| Mass (M⊙) | |||
| Radius (R⊙) | |||
| Luminosity (L⊙) | |||
| Density (cgs) | |||
| Surface gravity (cgs) | |||
| Effective Temperature (K) | |||
| Metallicity (dex) | |||
| Planetary Parameters: | TOI-1518b | ||
| Period (days) | |||
| Radius (RJup) | |||
| Mass (MJup) | |||
| Time of transit () | |||
| Optimal conjunction Time () | |||
| Semi-major axis (AU) | |||
| Inclination (degrees) | |||
| Equilibrium temperature (K) | |||
| Tidal circularization timescale (Gyr) | |||
| RV semi-amplitude (m/s) | |||
| Radius of planet in stellar radii | |||
| Semi-major axis in stellar radii | |||
| Transit depth (fraction) | |||
| Flux decrement at mid transit | |||
| Ingress/egress transit duration (days) | |||
| Total transit duration (days) | |||
| FWHM transit duration (days) | |||
| Transit Impact parameter | |||
| Blackbody eclipse depth at 2.5m (ppm) | |||
| Blackbody eclipse depth at 5.0m (ppm) | |||
| Blackbody eclipse depth at 7.5m (ppm) | |||
| Density (cgs) | |||
| Surface gravity | |||
| Safronov Number | |||
| Incident Flux (109 erg s-1 cm-2) | |||
| Time of eclipse () | |||
| Time of Ascending Node () | |||
| Time of Descending Node () | |||
| Mass ratio | |||
| Separation at mid transit | |||
| A priori non-grazing transit prob | |||
| A priori transit prob | |||
| Wavelength Parameters: | TESS | ||
| linear limb-darkening coeff | |||
| quadratic limb-darkening coeff | |||
| Telescope Parameters: | FIES | SOPHIE | |
| Relative RV Offset (m/s) | |||
| RV Jitter (m/s) | |||
| RV Jitter Variance | |||
| Transit Parameters: | TESS long cadence | TESS short cadence | |
| Added Variance | |||
| Baseline flux | |||
Appendix B Cross-Correlation-Function
Appendix C Global circulation models
Appendix D Retrieval analysis